OpenStudy (anonymous):

What is the equation of the line (in slope intercept form) that would be perpendicular to y= 1/5x + 1001 and go through the origin?

OpenStudy (anonymous):

John

OpenStudy (john_es):

The slope of the line the problem gives you is m=1/5. For the perpedicular line you have the formula, \[y=\frac{-1}{m}x+n\] where you have to find the value of n, knowing the perpendicular line pass through the origin. Then \[0=0+n\Rightarrow n=0\] So the perpendicular line will be, \[y=\frac{-1}{m}x=\frac{-1}{1/5}x=-5x\]

OpenStudy (anonymous):

can you please help me with this problem

OpenStudy (anonymous):

thank you

OpenStudy (john_es):

You're welcome.